A bag contains 4 red and 6 black balls. A sample of 4 balls is selected from the bag without replacement. Let X is the number of red balls. Find the p.d for X? a) Find probability of X greater than 2? b) Find probability of X for non-negative integers? 


1
Expert's answer
2021-05-31T00:50:34-0400

Let "X=" the number of red balls.


"P(X=0)=\dfrac{\dbinom{4}{0}\dbinom{6}{4}}{\dbinom{10}{4}}=\dfrac{1(15)}{210}=\dfrac{1}{14}"

"P(X=1)=\dfrac{\dbinom{4}{1}\dbinom{6}{3}}{\dbinom{10}{4}}=\dfrac{4(20)}{210}=\dfrac{8}{21}"

"P(X=2)=\dfrac{\dbinom{4}{2}\dbinom{6}{2}}{\dbinom{10}{4}}=\dfrac{6(15)}{210}=\dfrac{3}{7}"

"P(X=3)=\dfrac{\dbinom{4}{3}\dbinom{6}{1}}{\dbinom{10}{4}}=\dfrac{4(6)}{210}=\dfrac{4}{35}"

"P(X=4)=\dfrac{\dbinom{4}{4}\dbinom{6}{0}}{\dbinom{10}{4}}=\dfrac{1(1)}{210}=\dfrac{1}{210}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 & 4 \\ \hline 
 p(x) & 1/14 & 8/21 & 3/7 & 4/35 & 1/210 \\
\end{array}"

a)

"P(X>2)=P(X=3)+P(X=4)"

"=\dfrac{4}{35}+\dfrac{1}{210}=\dfrac{5}{42}"

b)


"P(X=0)=\dfrac{1}{14}"

"P(X=1)=\dfrac{8}{21}"

"P(X=2)=\dfrac{3}{7}"

"P(X=3)=\dfrac{4}{35}"

"P(X=4)=\dfrac{1}{210}"